One day, you are playing around with the change in your pocket. You decide to flip...

  1. One day, you are playing around with the change in your pocket. You decide to flip one coin and then you decide to flip two coins. Remember, a coin can land on either heads or tails and each coin flip is independent. Calculate the following probabilities:

One Coin Flip Probabilities:

P (heads) =

P (tails) =

Two Coin Flips Probabilities:

P (of at least one heads) =

P (of at least one tails) =

P (heads and heads) =

P (heads and tails) =

P (tails and tails) =

P (tails and heads) =
Answer #1

1). One Coin Flip Probabilities:

sample space :-{ H,T } ,  where H indicates a head, and T a tail.

total number of outcomes = 2

the needed probabilities be:-

P (heads) = \mathbf{\frac{1}{2}}
P (tails) = \mathbf{\frac{1}{2}}

2).Two Coin Flips Probabilities:

sample space :-{ HH , HT , TH , TT } ,  where H indicates a head, and T a tail.

total number of outcomes = 4

the needed probabilities be:-

answers calculation

P (of at least one heads) =

\mathbf{\frac{3}{4}}

the needed outcomes are: {HH,HT,TH}

number of outcomes = 3

probability = 3/4

P (of at least one tails) =

\mathbf{\frac{3}{4}}

the needed outcomes are: {TT,HT,TH}

number of outcomes = 3

probability = 3/4

P (heads and heads) =

\mathbf{\frac{1}{4}}

the needed outcome ={HH}

probability = 1/4

P (heads and tails) =

\mathbf{\frac{1}{4}}

the needed outcome ={HT}

probability = 1/4

P (tails and tails) =

\mathbf{\frac{1}{4}}

the needed outcome ={TT}

probability = 1/4

P (tails and heads) =

\mathbf{\frac{1}{4}}

the needed outcome ={TH}

probability = 1/4

*** if you have any doubt regarding the problem please write it in the comment box.if you are satisfied please give me a LIKE if possible..